OpenStudy (anonymous):

9-4z^2 would that be (3-2z)^2?

OpenStudy (anonymous):

no when you multiply (3-2z) * (3-2z) what is the result

OpenStudy (anonymous):

heh, true. 3-2z) (3+2z) ?

OpenStudy (anonymous):

you are right now.. its \[(3-2z) * (3+2z) = 3^2 - (2z)^2=9-4z^2\]